# Restless Bandit Problem with Rewards Generated by a Linear Gaussian Dynamical System

### v1

@@ 4.1 Regret Performance
Algorithm 1, OSS, has the following upper bound on regret (2). Proof is in Appendix C.3.
@@
Using Algorithm 1, regret (2) satisfies the following inequality with a probability of at least $(1-\delta)^{4}$:
$R_{n}\leq\sum_{a=1}^{k}2nB_{c}^{2}B_{R}^{2}\left(1-\left(1-\delta\right)^{4}\left(1-\exp\left(\frac{-4B\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)\right)\right),$ (16)
where $\Delta G_{c_{a}\mid\mathbf{c}}\triangleq G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}$, $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\triangleq\mathbb{E}\left[\Xi_{t}\left(\mathbf{c}\right)\Xi_{t}\left(\mathbf{c}\right)^{\top}\right]$, and $B\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)$ is a bound on uncertainty (13).
Theorem 2 proves that regret increases at worst linearly, i.e. $\mathcal{O}\left(n\right)$. First, the regret bound (16) decreases as the difference $\Delta G_{c_{a}\mid\mathbf{c}}$ or the reward covariance $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}$ increase. Second, the regret bound (16) decreases as uncertainty (13) decreases. The bound for $B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right)$ can be found in Appendix C.2.
Remark 3.
Although it is common in stationary bandits to see sublinear regret, linear regret may be the best possible rate of regret for non-stationary bandits (this is proven in [1]).
@@ 5 Numerical Results
where the process noise $\xi_{t}\sim\mathcal{N}\left(0,I_{4}\right)$ and the measurement noise $\eta_{t}\sim\mathcal{N}\left(0,1\right)$ are sampled from standard normal distributions. The LGDS (17) is proposed to study how the magnitude of the error covariance matrix $P_{\overline{a}}$ impacts performance of OSS, where $P_{\overline{a}}$ is directly impacted by the parameter $\theta\in[0,2\pi]$. Prior to the learner’s interaction with the LGDS (17), $10^{4}$ time steps are computed of the LGDS (17) to set the system to a steady state. After, the $10^{4}$ time steps, the length of the interaction between the environment and the learner is $n=10^{4}$ rounds. Regret (2) is used to provide a metric of performance. Parameter $s$ in algorithm 1, OSS, is set to $1$ in the top left plot. For comparison, we consider UCB [8], SW-UCB [9], and a learner that selects a random action each round (this learner is denoted as Random). We use UCB as an comparison since the eigenvalues of the LGDS (17) state matrix $\Gamma\in\mathbb{R}^{4\times 4}$ is Schur, implying that the reward distributions are stationary. SW-UCB is also used as a comparison since the reward is still generated by a dynamical system. Finally, Random is used as baseline for worst performance.
@@
Figure 2: Comparison algorithm’s regret normalized with respect to OSS’s regret. A positive percent implies that OSS has a lower regret than the compared algorithm.
@@
In the top left plot of Figure 2, regret for UCB (red), Sliding UCB (green) and Random (blue) is normalized with respect to OSS’s regret for each $\theta\in[0,2\pi]$. The middle plot of Figure 2 is the minimum eigenvalue of the Observability Gramian $\mathcal{O}$ [22] for both actions $c_{a}\in\{c_{1},c_{2}\}$, which is the solution of the Lyapunov equation $\mathcal{O}=\Gamma^{\top}\mathcal{O}\Gamma+c_{a}c_{a}^{\top}$. The bottom plot of Figure 2 is the real part of the eigenvalue of the state matrix $\Gamma$. In the white regions, all the comparison algorithms outperform OSS. Based on the plot in the middle, it appears that the low Observability Gramian minimum eigenvalue and a positive real part of the state matrix’s eigenvalue is the cause. For the blue regions, no algorithm outperforms Random, implying that the rewards are too noisy to estimate/predict for the compared algorithms. Finally, the gray regions is approximately where OSS performs the best, providing at least a 10% improvement for each of the algorithms mid-region. Based on the bottom 2 plots, this increase in performance is from the high observability and an eigenvalue with a negative real part for the state matrix. High observability lowers the magnitude of the error covariance matrix $P_{\overline{a}}$, which leads to a lower regret bound of OSS. In addition, an eigenvalue with a negative real part for the state matrix leads to rapid switching of the optimal action, making it difficult for UCB to adapt. For the plot on the far right, this is the relative performance of OSS for each parameter $s=1,2,3$ when the LGDS system (17) parameter set to approximately $85\pi/$ 8 (approximately where we see the largest improvement in performance of OSS in the top left plot). Therefore, it appears that as $s$ increases to $s=2,3$, regret performance of OSS decreases. In addition, regret for parameter $s=3$ is similar to regret of parameter $s=2$ as parameter $s=3$ may be identifying a more precise predictor but has more parameters to identify, which can cause a longer period of exploration.
@@ 6 Conclusion
We have presented an algorithm for addressing a variation of the restless bandit with a continuous state-space. The rewards generated by this restless bandit variation is a LGDS. Based on the formulation, we propose to learn a representation of the modified Kalman filter to predict the rewards for each action. We have shown that regardless of the sequence of actions chosen, the learned representation of the modified Kalman filter converges. It is then proven what strategy should be used given the bound on regret, leading to an optimism-based strategy.
In this work, we have not considered how the sequence of actions $c_{A_{t-s}},\dots,c_{A_{t-1}}$ impact prediction error and how to choose parameter $s$, i.e. how far the learner looks into the past. First, the perturbation added for exploration only considers error of the model and not the sequence of actions impact on the error of the prediction. In other words, the chosen sequence of actions are myopic. This is similar to what is studied in sensor selection problems where the sequence of chosen sensors impact the error of the Kalman filter state estimate. Therefore, future work will focus on the action sequence impact on the reward prediction. Next, an important parameter in OSS is the window size $s$, which is the number of rewards in the past that are used for predicting the reward for each action. In OSS, this is a parameter to set prior to the interaction with the environment. However, questions we care to ask is how to automate the process of choosing $s$.
@@ Appendix A Proof of Lemma 1
for any $c_{a}\in\mathcal{A}$ that is detectable with the LGDS (1). Since the LGDS (1) is detectable for any $c_{a}\in\mathcal{A}$, then the Kalman filter’s difference equation in (3) converges exponentially to its steady-state solution [24]. In addition, $P_{\overline{a}}\succeq P_{a}$ for any $c_{a}\in\mathcal{A}$. Therefore, $P_{\overline{a}}\succeq g\left(P_{\overline{a}},c_{a}\right)$ for any $c_{a}\in\mathcal{A}$. Let there be the modified Kalman filter (5) and the error covariance matrix $P_{t}^{\prime}\triangleq\mathbb{E}\left[\left(z_{t}-\hat{z}_{t}^{\prime}\right)\left(z_{t}-\hat{z}_{t}^{\prime}\right)^{\top}\mid\mathcal{F}_{t-1}\right]$. The iteration of $P_{t}^{\prime}$ is
$P_{t+1}^{\prime}=\left(\Gamma-\Gamma L_{a_{t}}c_{A_{t}}^{\top}\right)P_{t}^{\prime}\left(\Gamma-\Gamma L_{a_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{a_{t}}L_{a_{t}}^{\top}\Gamma^{\top}$ (18)
@@ Appendix B Proof of Theorem 1
Using the fact from Lemma 1 that
$g\left(P_{\overline{a}},c_{a}\right)\succeq g\left(P_{t}^{\prime},c_{a}\right),$
if $P_{\overline{a}}\succeq P_{t}^{\prime}$ for any $c_{a}\in\mathcal{A}$ and
$\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)P_{\overline{a}}\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{A_{t}}L_{A_{t}}^{\top}\Gamma^{\top}=g\left(P_{\overline{a}},c_{a}\right),$ (19)
then (18) has a steady-state solution for any $c_{A_{t}}\in\mathcal{A}$. Finally, based on Lyapunov method for (19), the matrix $\Gamma-\Gamma L_{a_{t}}c_{A_{t}}^{\top}$ is stable. In addition, there exists an action $c_{A_{t}}\in\mathcal{A}$ such that the error $\gamma_{A_{t}}\sim\mathcal{N}\left(0,c_{A_{t}}^{\top}P_{t}^{\prime}c_{A_{t}}+\sigma\right)$ has the following upper-bound
@@ C.1 Model Error
Let $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ and $G_{c_{a}\mid\mathbf{c}}$ be as in (10) and (8), respectively. Then, the matrix of the bias terms $\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}$ satisfies the following inequality with a probability of at least $1-\delta$ (the failure probability value $\delta$ is assumed to be in the open interval $(0,1)$):
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}.$ (22)
@@
where $\hat{Z}_{t}\triangleq\mathbb{E}\left[z_{t}z_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$ and $\hat{P}_{t}\triangleq\mathbb{E}\left[e_{t}e_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$. Since $\hat{Z}_{t}\succeq 0$ and $P_{t}\succeq 0$, then
@@
Therefore, using the Markov inequality [25], with a probability of at least $1-\delta$,
$\left|\beta_{A_{t}}\right|\leq\left\|c_{a}\right\|_{2}\frac{\sqrt{\mbox{tr}\left(\mathbb{E}\left[z_{t}z_{t}^{\top}\right]\right)}}{\delta}\leq\left\|c_{a}\right\|_{2}\frac{B_{R}}{\delta}\leq\frac{B_{c}B_{R}}{\delta}.$ (28)
@@
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}$ $\leq\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta}\left\|\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}$
$=\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$\leq\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)\cdot 1}$
$\overset{(*)}{\leq}\frac{B_{c}B_{R}}{\delta}\sqrt{\sum_{i=1}^{N_{a}}1}\sqrt{\sum_{i=1}^{N_{a}}\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$=\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\sum_{i=1}^{N_{a}}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$=\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\lambda I\right)\right)},$
$\Rightarrow\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)},$
where ()(*)( ∗ ) uses the Cauchy-Schwarz inequality. Since $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq\lambda I$ based on (11), then the difference $I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq 0$. Therefore, no further effort is needed to show that inequality (22) is satisfied with a probability of at least $1-\delta$.
@@
Let $\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ be identified based on (10) where the reward $\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ has expression (9). Then, the following inequality is satisfied with a probability of at least $\left(1-\delta\right)^{2}$, where $\delta\in(0,1)$ are designed parameters for tolerated failure probabilities:
$\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}\leq e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right),$ (29)
where $e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ and $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ are defined in (14) and (15), respectively.

### v2

@@ 4.1 Regret Performance
Algorithm 1, UBSS, has the following upper bound on regret (2). Proof is in Appendix C.3.
@@
Using Algorithm 1 and setting $\delta_{e}=\delta_{b}=\delta\in(0,1)$, regret (2) satisfies the following inequality with a probability of at least $(1-\delta)^{5}$:
$R_{n}\leq\sum_{t=1}^{s}\max_{c_{a}\in\mathcal{A}}\mathbb{E}\left[\left\langle c_{a^{*}}-c_{a},z_{t}\right\rangle\right]+\sum_{a=1}^{k}2(n-s)B_{c}^{2}B_{R}^{2}\left(1-\left(1-\delta\right)^{4}\left(1-\exp\left(\frac{-4B\left(\delta\mid\mathbf{c}\right)^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)\right)\right),$ (16)
where $\Delta G_{c_{a}\mid\mathbf{c}}\triangleq G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}$, $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\triangleq\mathbb{E}\left[\Xi_{t}\left(\mathbf{c}\right)\Xi_{t}\left(\mathbf{c}\right)^{\top}\right]$, and $B\left(\delta\mid\mathbf{c}\right)$ is
$B\left(\delta\mid\mathbf{c}\right)\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(s\lambda+(n-s)\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\lambda^{s/2}}\right)}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\\ +\sqrt{n-s}\frac{B_{c}B_{R}}{\delta}\sqrt{s}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}+\lambda B_{G}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}.$ (17)
Theorem 2 proves that regret increases at worst linearly, i.e. $\mathcal{O}\left(n\right)$. If the covariance for two different actions is large, e.g. large $\Delta G_{c_{a}\mid\mathbf{c}}$ and $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}$ terms, then the bound will decrease. The bound decreases exponentially as uncertainty (13) decreases.
@@ 5 Numerical Results
where the process noise $\xi_{t}\sim\mathcal{N}\left(0,I_{4}\right)$ and the measurement noise $\eta_{t}\sim\mathcal{N}\left(0,1\right)$ are sampled from standard normal distributions. The LGDS (18) is proposed to study how the magnitude of the error covariance matrix $P_{\overline{a}}$ impacts performance of UBSS, where $P_{\overline{a}}$ is directly impacted by the parameter $\theta\in[0,2\pi]$. Prior to the learner’s interaction with the LGDS (18), $10^{4}$ time steps are computed of the LGDS (18) to set the system to a steady state. After, the $10^{4}$ time steps, the length of the interaction between the environment and the learner is $n=10^{4}$ rounds. Regret (2) is used to provide a metric of performance. Parameter $s$ in Algorithm 1, UBSS, is set to $1$ in the top left plot. For comparison, we consider UCB [8], SW-UCB [9], and a learner that selects a random action each round (this learner is denoted as Random). We use UCB as a comparison since the eigenvalues of the LGDS (18) state matrix $\Gamma\in\mathbb{R}^{4\times 4}$ is Schur, implying that the reward distributions have a bounded covariance with a mean of zero. SW-UCB is also used as a comparison since the reward is still generated by a dynamical system. Finally, Random is used as baseline for worst performance.
@@
Figure 1: Comparison algorithm’s regret normalized with respect to UBSS’s regret. A positive percent implies that UBSS has a lower regret than the compared algorithm.
@@
In the top left plot of Figure 1, the percentage of UBSS’s regret (2) is lower than UCB (red), Sliding UCB (green) and Random (blue) regret is shown for each $\theta\in[0,2\pi]$. The middle plot of Figure 1 is the minimum eigenvalue of the Observability Gramian $\mathcal{O}$ [22] for both actions $c_{a}\in\{c_{1},c_{2}\}$, which is the solution of the Lyapunov equation $\mathcal{O}=\Gamma^{\top}\mathcal{O}\Gamma+c_{a}c_{a}^{\top}$. The bottom plot of Figure 1 is the real part of the eigenvalue of the state matrix $\Gamma$. In the white regions, all the comparison algorithms outperform UBSS. Based on the plot in the middle, it appears that the low Observability Gramian minimum eigenvalue and a positive real part of the state matrix’s eigenvalue is the cause. For the blue regions, no algorithm outperforms Random, implying that the rewards are too noisy to estimate/predict for the compared algorithms. Finally, the gray regions is approximately where UBSS performs the best, providing approximately a 10% improvement for each of the algorithms mid-region. Based on the bottom 2 plots, this increase in performance is from the high observability and an eigenvalue with a negative real part for the state matrix. High observability lowers the magnitude of the error covariance matrix $P_{\overline{a}}$, which leads to a lower regret bound of UBSS. In addition, an eigenvalue with a negative real part for the state matrix leads to rapid switching of the optimal action, making it difficult for UCB to adapt. For the plot on the far right, this is the relative performance of UBSS for each parameter $s=1,2,3$ when the LGDS system (18) parameter set to approximately $85\pi/$ 8 (approximately where we see the largest improvement in performance of UBSS in the top left plot). Therefore, it appears that as $s$ increases to $s=2,3$, regret performance of UBSS decreases. Since the number of parameters to identify increases exponentially as $s$ increases (leading to longer exploration times), regret performance of UBSS decreases as $s$ increases.
@@ 6 Conclusion
We have presented an algorithm for addressing a variation of the restless bandit with a continuous state-space. The rewards generated by this restless bandit variation is a LGDS. Based on the formulation, we propose to learn a representation of the modified Kalman filter to predict the rewards for each action. We have shown that regardless of the sequence of actions chosen, the learned representation of the modified Kalman filter converges. It is then proven what strategy should be used given the bound on regret, leading to an uncertainty-based strategy.
In this work, we have not considered how the sequence of actions impact prediction error, how to choose window size (how far the learner looks into the past), and best obtainable performance of SMAB with LGDS environments. First, the perturbation added for exploration only considers error of the model and not the sequence of actions impact on the error of the prediction. In other words, the chosen sequence of actions are myopic. Therefore, future work will focus on the action sequence impact on the reward prediction. Next, an important parameter in UBSS is the window size. In UBSS, this is a parameter to set prior to the interaction with the environment. However, questions we care to ask is how to automate the process of choosing window size. Finally, UBSS has linear regret performance. Therefore, future work will be to derive the best obtainable performance of any algorithm applied to a SMAB with rewards generated by this paper’s proposed LGDS. We will then analyze if UBSS regret performance is close or far to the best obtainable performance.
@@ Appendix A Proof of Lemma 1
for any $c_{a}\in\mathcal{A}$ that is detectable with the LGDS (1). Assuming that $P_{\overline{a}}\succeq P$, we satisfy the following inequalities according to [23]:
$P_{\overline{a}}\overset{(*)}{\succeq}g\left(P_{\overline{a}},c_{A_{t}}\right)$ $=\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)P_{\overline{a}}\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{A_{t}}L_{A_{t}}^{\top}\Gamma^{\top}$
$\succeq\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)P\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{A_{t}}L_{A_{t}}^{\top}\Gamma^{\top}.$
For ()(*)( ∗ ), assume for proof of contradiction that $g\left(P_{\overline{a}},c_{A_{t}}\right)\succeq P_{\overline{a}}$. Based on [23], then $g\left(g\left(P_{\overline{a}},c_{A_{t}}\right),c_{A_{t}}\right)\succeq g\left(P_{\overline{a}},c_{A_{t}}\right)$, where the iteration will either diverge to infinity or $P_{A_{t}}$. However, we assumed that $P_{\overline{a}}\succeq P_{A_{t}}$ and the steady-state solution $P_{A_{t}}=g\left(P_{A_{t}},c_{A_{t}}\right)$ exists from Assumption 3. Since there is a contradiction, ()(*)( ∗ ) is true. Therefore, based on above, let there be the modified Kalman filter (5) and the error covariance matrix $P_{t}^{\prime}\triangleq\mathbb{E}\left[\left(z_{t}-\hat{z}_{t}^{\prime}\right)\left(z_{t}-\hat{z}_{t}^{\prime}\right)^{\top}\mid\mathcal{F}_{t-1}\right]$. The iteration of $P_{t}^{\prime}$ is
$P_{t+1}^{\prime}=\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)P_{t}^{\prime}\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{A_{t}}L_{A_{t}}^{\top}\Gamma^{\top},$ (19)
@@ Appendix B Proof of Theorem 1
The following is true based on Lemma 1 and Assumption 3:
  • The matrix pair $\left(\Gamma,Q^{1/2}\right)$ is controllable.
  • The matrix $\sigma^{2}\Gamma L_{A_{t}}L_{A_{t}}^{\top}\Gamma^{\top}\succeq 0$.
  • The iteration (19) is bounded by $P_{\overline{a}}$ and is monotonic.
Therefore, the matrix $\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}$ is stable. In addition, there exists an action $c_{A_{t}}\in\mathcal{A}$ such that the error $\gamma_{A_{t}}\sim\mathcal{N}\left(0,c_{A_{t}}^{\top}P_{t}^{\prime}c_{A_{t}}+\sigma\right)$ has the following upper-bound
@@ C.1 Model Error
Let $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ and $G_{c_{a}\mid\mathbf{c}}$ be as in (10) and (8), respectively. Then, the matrix of the bias terms $\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}$ satisfies the following inequality with a probability of at least $1-\delta_{b}$ (the failure probability value $\delta_{b}$ is assumed to be in the open interval $(0,1)$):
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}.$ (22)
@@
where $\hat{Z}_{t}\triangleq\mathbb{E}\left[z_{t}z_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$ and $P_{t}\triangleq\mathbb{E}\left[e_{t}e_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$. Since $\hat{Z}_{t}\succeq 0$ and $P_{t}\succeq 0$, then
@@
Therefore, using the Markov inequality [24], with a probability of at least $1-\delta_{b}$,
$\left|\beta_{A_{t}}\right|\leq\left\|c_{a}\right\|_{2}\frac{\sqrt{\mbox{tr}\left(\mathbb{E}\left[z_{t}z_{t}^{\top}\right]\right)}}{\delta_{b}}\leq\left\|c_{a}\right\|_{2}\frac{B_{R}}{\delta_{b}}\leq\frac{B_{c}B_{R}}{\delta_{b}}.$ (28)
@@
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}$ $\leq\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\left\|\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}$
$=\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$\leq\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)\cdot 1}$
$\overset{(*)}{\leq}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\sum_{i=1}^{N_{a}}1}\sqrt{\sum_{i=1}^{N_{a}}\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$=\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\sum_{i=1}^{N_{a}}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$=\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\lambda I\right)\right)},$
$\Rightarrow\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)},$
where ()(*)( ∗ ) uses the Cauchy-Schwarz inequality. Since $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq\lambda I$ based on (11), then the difference $I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq 0$. Therefore, no further effort is needed to show that inequality (22) is satisfied with a probability of at least $1-\delta_{b}$.
@@
Let $\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ be identified based on (10) where the reward $\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ has expression (9). Then, the following inequality is satisfied with a probability of at least $\left(1-\delta_{e}\right)\left(1-\delta_{b}\right)$, where $\delta_{e},\delta_{b}\in(0,1)$ are designed parameters for tolerated failure probabilities:
$\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}\leq e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right),$ (29)
where $e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)$ and $b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)$ are defined in (14) and (15), respectively.